23.12 grams of glucose (C6H12O6) is dissolved in 1.00 kg of water. Calculate the molality

1
Expert's answer
2022-05-18T22:45:02-0400

1. we find the amount of matter (mol) of glucose. To do this, we divide the given mass of glucose (180g/mol) by its molecular mass:

n = m / M = 23.12 / 180 = 0.1284 mol.

2. we find the molal concentration. To do this, we divide the amount of substance (n, mol) of glucose by the mass of the solvent, is water (kg):

Cmolality = 0.1284 mol / 1 kg = 0.1284

ANSWER: molality=0.1284